OpenStudy (lilai3):

Meteorites Problem An astronomer found 3 meteorites weighing 9.4 pounds, 5.7 pounds, and 24.5 pounds. If 1 kilogram is about 2.2 pounds, find the average mass of the meteorites in kilograms.

OpenStudy (anonymous):

9.4/2.2lbs = 4.27 kg 5.7/2.2lbs = 2.59kg 24.5/2.2lb = 11.13kg add all 3 and you get 17.99 , pretty much 18 18/3 = 6 kg

OpenStudy (lilai3):

thanks, @Pandae

OpenStudy (lilai3):

and thanks for your time, experimentx!

OpenStudy (experimentx):

sure ...